OpenStudy (anonymous):

Question about average velocity: I can understand why 'average velocity' equals change of distance/ change of time. mathematical equation: average velocity= average distance/average time. v=d/t. What I don't get is "Why is 'average velocity' also equal to initial velocity + final velocity over 2?" mathematical equation: average velocity = (Vi + Vf)/(2).

OpenStudy (lastdaywork):

Do you want to know where did the "...mathematical equation: average velocity = (Vi + Vf)/(2)..." came from? Or are you looking for a physical interpretation of average velocity ??

OpenStudy (anonymous):

I am looking for the 1st question you asked me. I get the physical interpretation of average velocity. Your help would be great.

OpenStudy (lastdaywork):

"...mathematical equation: average velocity = (Vi + Vf)/(2)..." is applicable only when acceleration is constant. For proof, consider a v-t graph with constant acceleration (hence straight line). Calculate average velocity using "...'average velocity' equals change of distance/ change of time..." Note that the total displacement would be the area under the graph. I leave the rest as an exercise. :)

OpenStudy (lastdaywork):

Consider this as barter system :) Please explain the meaning of "정현돈" in English.

OpenStudy (anonymous):

Thank you for your answer. ^^ I am not quite sure of it's meaning. The characters are in Korean, and its meaning is derived from Hanja.

OpenStudy (vincent-lyon.fr):

What LastDayWork wrote: "...mathematical equation: average velocity = (Vi + Vf)/(2)..." is applicable only when acceleration is constant." is very important. In no case is this equation a universal one. It is only a useful trick to express average velocity and distance travelled during a given time interval when acceleration is constant. What I want to say is that this equation is useful if you have to solve small problems with multiple choice answers in the context of a competition where you have to be as fast as possible. On the other hand, it is of no use for your true understanding of the concepts in mechanics.

OpenStudy (lastdaywork):

@정현돈 Please consider closing the question.

OpenStudy (anonymous):

Thank you for a good insight, Mr. Lyon.Fr. Your explanation excels for my understanding and my studies.